Circular No. 7361

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVA 2000H IN IC 454 Independent discoveries of an apparent supernova in IC 454 have been

reported by M. Armstrong, Rolvenden, Kent, England, on unfiltered CCD images taken on Feb. 7.97 and 8.03 UT (mag ∼ 17) for the U.K. Nova/Supernova Patrol with a 0.30-m telescope, and by M. Papenkova and W. D. Li, University of California at Berkeley, on behalf of the Lick Observatory Supernova Search (cf. IAUC 6627, 7126), on unfiltered CCD images taken on Feb. 8.18 and 8.22 (mag ∼ 17.3) with the 0.8-m Katzman Automatic Imaging Telescope (KAIT). KAIT measurements show that the new object is located at α = 6^{h}51^{m}07^{s}.67, δ = +12^{o}55′18′′.5 (equinox 2000.0), which is 19′′.6 east and 0′′.2 south of the nucleus of IC 454. Armstrong provides the following position end figures from his exposure on Feb. 8.782: 07^{s}.60, 18′′.4 (or 25′′ east and 2′′ north of the galaxy center). G. M. Hurst, Basingstoke, England, notes that confirming images taken by S. Foulkes, Ashperton Village, on Feb. 8.812 and by T. Boles, Wellingborough, on Feb. 9.11 show SN 2000H near mag 17. SN 2000H also appears on earlier KAIT images taken on Jan. 27.2 (mag ∼ 19.0) and Feb. 2.2 (mag ∼ 17.3). Earlier patrol images by Armstrong appear to show the new object on Jan. 23.016 (trace) and 25.938 (mag 18.5; limiting mag 19.0). SN 2000H does not appear on CCD frames taken on 1999 Dec. 5 (limiting mag 19.0) and 2000 Jan. 4/5 by Armstrong, or on a KAIT image taken on Jan. 8.2 (limiting mag ∼ 20.0). Armstrong adds that nothing is visible at the location of SN 2000H on the second Palomar Sky Survey image.

AX J0043–737 J. Yokogawa and K. Koyama, Kyoto University, write: “An ASCA ob-

servation of the Small Magellanic Cloud (SMC) on 1999 May 10–11 detected coherent pulsations of period 87.58073 ± 0.00004 ms with a significance of 99.98 percent from AX J0043–737, which is located at α = 0^{h}42^{m}35^{s}, δ = −73^{o}40′30′′ (equinox 2000.0; ± 1′ at 90-percent confidence). The spectrum was described by a power law with photon index 1.7_{−0.5}^{+0.9}. The x-ray flux in the band 0.7–10.0 keV was 2.0 × 10^{−13} erg s^{−1} cm^{−2}, corresponding to a luminosity of 8.6 × 10^{34} erg/s for a distance of 60 kpc. We suggest that AX J0043–737 is a Crab-like pulsar in the SMC, although the possibility of an x-ray binary with a short pulse period is not excluded. Within the ASCA error region, a ROSAT source RX J0042.6–7340 is found (Kahabka et al. 1999, A.Ap. Suppl. 136, 81). Confirmation attempts at x-ray, radio, and other wavelengths are encouraged.”

2000 February 9  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2000 CBAT  Daniel W. E. Green